Algebraically find the x and y intercepts of the equation 3x - 5y = 10 and write each as an ordered pair.

Answer by Mathtut(3670)

the y intercept occurs when x=0 and the x intercept occurs when y=0 so lets take the y intercept first 3(0)-5y=10 so -5y=10 y=-2 so the y intercept is -2 and written as an ordered pair would be (0,-2) now lets handle the x intercept remembering that y=0 so we restate our equation this time replacing the y value with 0 so we get 3x-5(0)=10 so 3x=10 so x=10/3 so x-intercept is at 10/3 written as an ordered pair would be (10/3,0)